The odds against student A solving a business statistics problems are 4:3 and odds in favor of student B solving the same problem are 7:8.

i) what is the chance that the problem will be solved if they both try?

ii) What is the probability that neither solves the problem?


1
Expert's answer
2021-03-01T07:11:55-0500

From conditions we suppose that probabilities of successful solved problems are equal to 3/7 and 7/15 for A and B students accordingly.

b) First let's find the probability that neither solves the problem. It's equal to (1 - 3/7) * (1- 7/15) = 4/7 * 8/15 = 32 / 105, where we multiply probabilities of unsuccessful outcomes.

a) Using the probability from the previous point, we can find the probability that the problem will be solved if they both try. It's equal to (1 - 32/105) = 73 / 105, because it's enough for success if at least one of them solve the problem.